ok here is the question and i am completely dumbfounded about these types of questions. here is the question "write an equation in slope-intercept form of the line passing through each pair of points. (3,5) and (2,1).. thnks i will be greatful if you can help.
Answer by Alan3354(69443)

"write an equation in slope-intercept form of the line passing through each pair of points. (3,5) and (2,1)
----------------
You're missing a "
------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.